The Sum of three numbers is 123. The Second Number is 9 Less than two times the first number. The third number is 6 more than three times the first number. Find the three numbers
Answer by checkley77(12569)

x+y+z=123
y=2x-9
z=3x+6
x+(2x-9)+(3x+6)=123
6x-3=123
6x=123+3
6x=126
x=126/6
x=21 for the first number.
2*21-9=42-9-33 for the second number.
3*21+6=63+6=69 for the third number.
proof:
21+33+69=123
123=123